OpenStudy (anonymous):

Solve 2sin^2 x -5sinx-3=0

OpenStudy (anonymous):

Call sin x = x. This is a quadratic equations: 2x^2 - 5x -3 = 0. Solve by factoring: (2x + 1)((x - 3) = 0. a) x = sin x = -1/2 = sin 7Pi/6 and sin 11Pi/6 b) x = sin x = 3: rejected as > 1 x = .... and x = ..... (0 < x < 2Pi)

OpenStudy (anonymous):

How do these threads that are 13 hours old suddenly come up in the menu on the left???

OpenStudy (anonymous):

bc i bumped it

OpenStudy (anonymous):

ahhh! i see...